### v2

# Subhalos are Anisotropically Distributed and Aligned with the Smooth Matter Distribution of Their Host Halos

Abstract
We investigate the distributions of subhalos about their hosts in two suites of zoom-in N-body simulations of halo growth – one suite focused on Milky Way-Mass halos ($\sim 10^{12}\mathrm{M}_{\odot}$) and another focused on cluster-mass halos ($\sim 10^{15}\mathrm{M}_{\odot}$) in the Symphony simulation suite. We find, in agreement with previous work on this subject, that subhalos are distributed anisotropically about their host halos. In particular, the positions of subhalos lie preferentially near the major axes of their host halos, possibly implying that satellite galaxies will exhibit a similar alignment. Furthermore, we show that in two-dimensional projection subhalos are more likely to be observed near the halo center (where the central galaxy presumably resides) when the host halo is projected nearly along its major axis. This projection effect is significant. Within projected radii of a few percent of the virial radius of the host halo, the fraction of mass in subhalos is $\sim 175\%$ larger for Milky Way mass halos and as much as $\sim 195\%$ larger for cluster halos when projected along the major axis as compared to the average from a random projection. This result has consequences for many applications including the interpretation of gravitational lenses. Finally, we find that the orbital angular momentum vector of subhalos is aligned with the angular momentum vector of their host halo, indicating that a significant component of a halo’s angular momentum may be carried in its subhalos. This has consequences for galaxy formation models which use host halo angular momentum as a proxy for galaxy momentum.
keywords:
galaxies: haloes – large-scale structure of Universe – cosmology: dark matter – galaxies: clusters: general – gravitational lensing: strong
pubyear: 2022pagerange: Subhalos are Anisotropically Distributed and Aligned with the Smooth Matter Distribution of Their Host HalosReferences
1 Introduction
[…]
One tool for detecting dark matter substructure is strong gravitational lensing – a phenomena which takes place when a massive foreground object, such as a dark matter halo and the galaxy that it contains, deflects the light of a background luminous object, such as a galaxy or quasar. For strong lensing to occur, the two-dimensional projected density of the “lens” must be greater than a critical density. This threshold is often only reached in the area near the centers of halos with size defined by the Einstein radius. For substructure lensing to occur, in particular, the projected positions of subhalos must be near the scale of the Einstein radius. For example, lensing by substructure has been noted in multiply-imaged systems (Meneghetti et al., 2006; Alard, 2008; Hezaveh et al., 2016) as well as arc systems (Mao & Schneider, 1998; Metcalf & Madau, 2001; Bradač et al., 2002; Metcalf & Zhao, 2002; Chiba & Takahashi, 2002; Dalal & Kochanek, 2002; Keeton, 2003; Metcalf et al., 2004; Oguri et al., 2012). If subhalos are distributed anisotropically with respect to their host halos, it is more likely that subhalos will fall in close proximity to the Einstein radius when a halo is viewed at a specific orientation. In other words, the probability of lensing by substructure is not uniform, but instead is biased by halo shape and halo orientation. The goal of this work is to explore the anisotropic distribution of subhalos about their hosts halos as a prelude to forthcoming studies of strong gravitational lensing. However, the anisotropic distributions of subhalos (and, by assumption, of satellite galaxies) about their hosts can have consequences for many kinds of observations, some of which we discuss briefly below.
A significant body of work predicts that subhalos in CDM simulations of structure formation should be distibuted anisotropically about their host halos. Host dark matter halos are ellipsoidal in shape (Barnes & Efstathiou, 1987; Dubinski & Carlberg, 1991; Warren et al., 1992; Dubinski, 1994; Tormen, 1997; Thomas et al., 1998; Jing & Suto, 2002; Springel et al., 2004; Hopkins et al., 2005), and in particular, halos tend to be prolate ellipsoids. Zentner et al. 2005b found preferential alignments between the overall mass distribution of Milky Way-mass host halos and the spatial distribution of their subhalos, and this was confirmed in galaxy formation simulations shortly thereafter (Libeskind et al., 2007). More recently, Karp et al. (2023) find an overdensity of satellite galaxies along the major axis of their central galaxy in the IllustrisTNG simulation. It is believed that these phenomena are the result of halos preferentially merging along filaments in the large-scale structure (Wang et al., 2005; Zentner et al., 2005b; Libeskind et al., 2007, 2011, 2015; Shi et al., 2015; Kang & Wang, 2015; Shao et al., 2018; Morinaga & Ishiyama, 2020). These filaments trace out preferred directions about halos in the large scale structure. Several simulations have found that alignments between neighboring dark matter halos tend to take place along the direction of the longest axes (the major axes) of the halos. Faltenbacher et al. 2002 and Hopkins et al. 2005 detect alignments between the major axes of cluster-sized halos out to separations of approximately 100 $\textrm{h}^{-1}$ Mpc and 30 $\textrm{h}^{-1}$ Mpc, respectively. Using the IllustrisTNG simulation, Rodriguez et al. (2024) reports alignments out to 10 $\textrm{h}^{-1}$ Mpc for both the stellar and dark matter components of galaxies.
The prediction that subhalos are disributed anisotropically about their hosts and that they are aligned preferentially with the non-spherical dark matter distribution of their hosts has some observational support from a variety of analyses of galaxies from surveys such as the Sloan Digital Sky Survey (SDSS). Using a sample of galaxy groups from SDSS DR4, Brainerd (2005) and Yang et al. (2006) showed that satellites tend to be aligned with the major axis of the central galaxy in the group. Azzaro et al. (2007) reached qualitatively similar conclusions for the SDSS DR4 sample using complementary methods (see also Azzaro et al., 2006). Wang et al. (2018) and Rodriguez et al. (2022) performed a similar analyses using SDSS DR7 and SDSS DR16 data respectively and found that this alignment exists primarily between red central galaxies and their red satellites. More recently, Samuels & Brainerd (2023) showed that satellite distributions tend to be lopsided with satellites preferentially residing on one side of their host in agreement with Libeskind et al. (2016) and Brainerd & Samuels (2021). Similar results have been observed in cluster lensing analyses. Using SDSS Evans & Bridle (2009), determined that dark matter halos are preferentially aligned in the same direction as indicated by the spatial distribution of cluster galaxies using cluster lensing. Gonzalez et al. (2021) drew similar conclusions from a galaxy lensing analysis of a number of surveys.
In contrast to the large-scale structure survey analyses mentioned above, satellites of the Milky Way and M31 galaxies have been observed to preferentially lie in a plane perpendicular to their host galaxy major axis (Lynden-Bell, 1976; Kunkel & Demers, 1976; Majewski, 1994; Mateo, 1998; Grebel, 1998; Hartwick, 2000; Willman et al., 2004; Kroupa et al., 2005; Metz et al., 2009; Pawlowski et al., 2012; Conn et al., 2013; Ibata et al., 2013). This particular alignment of satellites has been dubbed the Holmberg Effect and was first described in Holmberg (1969). Aside from the fact that these Local Group measurements refer to only two central galaxies, there are other differences between the Holmberg Effect in the Local Group and the SDSS measurements. First, and foremost, the central galaxies of the Local Group are both disk galaxies, whereas the statistics of the SDSS measurements were dominated by elliptical galaxies. Second, the Milky Way and M31 form a close pair of galaxies and it is uncertain the degree to which alignment effects may be enhanced or reduced in close pair systems.
In addition to the biases in the angular distributions of subhalos about their host galaxies, subhalos may also be kinematically biased relative to the smooth dark matter component in their hosts. These biases may be relevant in studies of galaxy formation. In models of disk galaxy formation, the central disk galaxy acquires its angular momentum from its host halo (Navarro & Benz, 1991; Navarro & White, 1993; Navarro & Steinmetz, 1997; D’Onghia et al., 2006; Kaufmann et al., 2007). For this reason, host halo angular momentum is often used as a proxy for galaxy angular momentum in semi-analytic galaxy formation models (Somerville et al., 2008; Benson, 2012). However, the influence of subhalos on angular momentum must not be overlooked in these models. Subhalos that merge from coherent directions, such as along filaments, often have angular momenta that are aligned with the angular momentum of their host halo (Aubert et al., 2004). However, it is not clear that the angular momentum carried by the subhalos should be included in the available reservoir of angular momentum that can be conveyed to the central disk galaxy. If it is the case that subhalos carry more angular momentum per mass than hosts in general and that subhalo angular momentum displays some preferred alignment with the host halo, this could induce a bias in models that estimate the amount of halo angular momentum transmitted to the baryonic disk. Whether or not the subhalos represent a distinct population and whether or not they represent the angular momentum of typical host particles should be accounted for in such models.
In this paper, we revisit the question of the degree to which the positions and momentum of subhalos within host halos are aligned with the mass distribution of their hosts, particularly, whether or not subhalos are preferentially aligned with the principle axes and angular momenta of their hosts. As will be evident, the primary application we have in mind is strong gravitational lensing, though we have attempted to summarize results in such a way as to be useful for a variety of other applications. This paper is organized as follows. In Section 2 of this paper we discuss the simulations used, Section 3 introduces the various halo properties that we studied, and in Section 4 and Section 5 we go over the alignment results for host halos and host-sub halo alignments, respectively. Finally we discuss our results and draw conclusions in Section 6.
2 Simulations
In this section, we briefly discuss the simulation data as well as the procedure that we use to construct our various data sets. The two simulations we describe are part of the Symphony suite of cosmological zoom-in simulations (Nadler et al., 2023). We refer the reader to Nadler et al. (2023) for a more detailed description of the simulations. The technique that we use for extracting data sets of host halos and subhalos is the same as that used in both Mezini et al. (2023) and Fielder et al. (2020). We refer the reader to these earlier works for more detailed descriptions of this procedure.
2.1 Milky Way and Cluster Mass Zoom-In Simulations
[…]
2.2 Halo Identification
Dark matter halos in these simulations were identified by the $\tt{ROCKSTAR}$ (version 0.99.9-RC3+) halo finder (Behroozi et al., 2013). In summary, $\tt{ROCKSTAR}$ is a 6D phase-space-based finder that produces catalogs of all the halos identified as well as tables of all the particles associated with each halo. A more detailed description of $\tt{ROCKSTAR}$ can be found in Behroozi et al. (2013); the source code is publicly available [^1] https://bitbucket.org/gfcstanford/rockstar. We use the tables of particles which are produced for each halo to analyze subhalo positions and angular momenta relative to that of their hosts.
Our $\tt{ROCKSTAR}$ based definition of halo components is the same as in the previous works of Fielder et al. (2020) and Mezini et al. (2023) and translates into the following three groupings of halo particles.
  • host only: These are particles that are associated with the host halo but not associated with any subhalos identified by $\tt{ROCKSTAR}$. The host only particles are part of the smooth component of the host halo or diffuse structures such as streams that are not identified by $\tt{ROCKSTAR}$.
  • subhalo only: These are particles associated with at least one subhalo within the host halo, as identified by $\tt{ROCKSTAR}$. Subhalos within the symphony simulation are defined such that they contain a minimum of 300 particles at z=0.
  • total halo: These are particles associated with the host halo, including all particles associated with subhalos within the host. In the vast majority of analyses in the literature, this is the set of particles that is used to characterize a halo; however, both Fielder et al. (2020) and Mezini et al. (2023) have found that the smooth, host only component of the host halo is markedly different from the total halo particles in a number of important ways.
From the preceding definitions, it follows that combining the particles in the host only and subhalo only samples recover the set of particles in the total halo particle set.
3 Quantifying Halo Shape and Orientation
3.1 The Major Axis of the Host Halo
The shapes of dark matter halos are most often characterized by the ratios of the lengths of their principle axes. We will refer to the principle semi-axes lengths as $a$, $b$, and $c$, so that $a>b>c$. The major axis length is thus $a$. Dark matter halos are most often found to be nearly prolate ellipsoids, which have $a>b\sim c$, but they can also be triaxial ellipsoids where $a>b>c$, and $b$ and $c$ are significantly different from one another.
The lengths of the principle axes can be computed as the eigenvalues of the inertia tensor, while their orientations are specified by the eigenvectors. The largest eigenvalue specifies $a$, while the smallest eigenvalue is $c$. The inertia tensor is
$I_{ij}=\sum_{n}x_{i,n}x_{j,n},$ (1)
where $x_{i,n}$ and $x_{j,n}$ are the coordinates of the $n^{\rm th}$ particle and the sum is over all particles identified with a halo. All particles in our simulations have identical masses, so there is no need to include the particle masses in the sum above.
[…]
We find the directions of the principle axes of the halo by calculating the inertia tensor and its eigenvectors directly from the particle data. We find the inertia tensor for the subhalo only component of each halo by subtracting the inertia tensor of the host only component from that of the total halo set. We then diagonalize this inertia tensor to get the subhalo component principle axes.
Figure 1: Distributions of halo shape as measured by the ratio of the halo minor axis to major axis ($c/a$) and the semi-major to major axis ($b/a$). The filled histograms correspond to c/a and hollow histograms to $b/a$. Milky Way-mass halos are shown in panel ($a$) on the left and cluster-mass halos on right in panel ($b$). Cluster-mass halos are generally less spherical.
3.2 Halo and Subhalo Angular Momentum
Similar to the major axis, angular momentum is also calculated directly from particle data. Per particle, we calculate the cross product of $\vec{r}_{n}$, the position of the $n^{\rm th}$ particle with respect to the host halo’s center, and $\vec{v}_{n}$, the velocity of $n^{\rm th}$ particle, also with respect to the host halo center. Summing over the angular momenta of all particles gives the total halo angular momentum,
$\vec{J}=\sum_{n}m_{n}(\vec{r}_{n}\times\vec{v}_{n}),$ (2)
where $m_{n}$ is the mass of the $n^{\rm th}$ particle. The total angular momentum, $\vec{J}_{\mathrm{Tot}}$, includes the angular momentum associated with both host halo and subhalo particles and is computed by summing over the total halo particle set. The host angular momentum, $\vec{J}_{\mathrm{Host}}$, is calculated using Equation 2 by summing over the host only particle set, while the angular momentum carried by the subhalos, $\vec{J}_{\mathrm{Sub}}$, is defined to be the difference $\vec{J}_{\mathrm{Sub}}$ = $\vec{J}_{\mathrm{Tot}}$ - $\vec{J}_{\mathrm{Host}}$.
3.3 The Subhalo Mass Distribution
We are interested in the way in which the substructure of a halo is distributed about its host. The distribution of subhalos around their hosts in two-dimensional projection and whether this distribution varies based on the projection angle are of special interest due to their relevance to gravitational lensing. Two-dimensional projections are an interesting special case because the projected 2D density is the relevant quantity in lensing which results in the angular deflections that distort the source. A number of strong lensing analysis programs count among their goals to infer the properties of subhalos (and other line-of-sight halos) in projection (e.g., Oguri & Marshall, 2010; Mckean et al., 2015; Ivezic et al., 2019). To this end, we calculate the projected mass in subhalos near the halo center from different viewing angles, including projections along the major axis, for which the projected mass density is highest. We recommend referring to Fig. 2 while reading the steps of this calculation. Our procedure is as follows.
  1. 1.
    Define a set of axes along which to project host halo and subhalo mass. The axes are centered on the host halo and spaced isotropically over the surface of a sphere.
  2. 2.
    Define a cylinder, with a central axis that points along the halo major axis and intersects the halo center such that the circular face of the cylinder has a radius $r<r_{\rm{VIR}}$ and the length of the cylinder extends to the virial radius of the halo.
  3. 3.
    Measure the total mass of subhalo particles that fall within the cylinder and divide it by the total halo mass to calculate the subhalo mass fraction within the cylinder.
  4. 4.
    Rotate the cylinder about the halo center so that it points along the subsequent axis in the predefined set of projection axes.
  5. 5.
    Calculate the angular separation, $\theta$, between the major axis and the new axis of projection.
  6. 6.
    Repeat the mass measurement from step (iii).
  7. 7.
    Repeat the process of rotating and measuring the subhalo mass [steps (iii)-(vi)] within each rotated cylinder for all the axes in the predefined set of axis directions from step (i).
Figure 2: Diagram showing how cylinders are defined to calculate projected mass as described in Section $3.$ The dotted pink circle represents the region enclosed within the virial radius of the halo. The light pink ellipsoid represents the shape of the host halo even though all particles within the virial radius are used, and the filled in dark pink circles represent subhalos. The left image shows a cylinder of radius $r$ oriented near the major axis. The cylinder is then rotated to some random orientation as shown in the right image. When comparing the left and right images, we can see how the number of subhalos within cylinders of the same size can vary depending on orientation.
This procedure is repeated for each halo of the Milky Way-mass and cluster-mass halos separately for cylinders with logorithmically-spaced radii from 0.001 to 1 times their host virial radius. By definition, a radius of $r=0$ encompasses none of the halo mass, whereas a cylinder of radius $r=r_{\rm{VIR}}$ encompasses the entire halo. Once completed, for both cluster-mass and Milky Way-mass halos, we will have a distribution of subhalo mass fractions within a cylinder of radius $r$ as a function of the angle $\theta$ between the cylinder’s axis and the halo’s major axis.
4 Results I: Host Halo Alignments
We begin our discussion of halo alignments with a study of the properties of particles that are associated with the host halo. It has been found that the orientation of the halo angular momentum vector with respect to the principle axes of the halo is not isotropic. Rather, the angular momentum vector is correlated with the minor axis of the halo (e.g., Warren et al., 1992; Tormen, 1997), as expected from tidal torque theory (e.g., Porciani et al., 2002a, b). The halo major and minor axes are orthogonal to one another (by definition). Therefore, this suggests that the angular momentum of a halo should be oriented in a manner that is preferentially orthogonal to its major axis.
In Fig. 3 we show the cumulative probability distribution of halos as a function of the cosine of the angular separation, $|\cos(\theta)|$, where $\theta$ is the angle between the host halo angular momentum vector, $\vec{J}_{\mathrm{Host}}$, and the host major axis. The angle $\theta$ plays the role of a zenith angle with zero along the angular momentum vector, so a model in which angular momenta and major axes are uncorrelated would yield a uniform distribution of $|\cos(\theta)|$. Panel (a) corresponds to Milky Way-mass halos and (b) to cluster-mass. In each panel, the solid lines represent the cumulative distribution function of alignments from our simulation data sets. The straight dotted lines represent the distribution expected for a random, isotropic relation between angular momentum and major axis. The filled areas represent errors relative to an isotropic underlying true distribution. These errors were estimated by building 1000 mock samples of alignments from an isotropic underlying distribution. Each mock sample contained the same number of data points as the simulation data (45 for Milky Way mass and 95 for cluster mass). The errors are then the $\rm{16^{th}}$ and $\rm{84^{th}}$ percentiles of these mock samples.
As discussed above, we expect that the orientation of $\vec{J}_{\mathrm{Host}}$ with respect to host halo major axes is not isotropic and will be preferentially aligned with the minor axes, and perpendicular to the major axes. Our data is suggestive of an orthogonal alignment between the major axis and angular momentum vector, $\vec{J}_{\mathrm{Host}}$, as expected. In both panels of Fig. 3, we see an excess of halos for which the angle between the angular momentum vector and the major axis, $|\cos(\theta)|$, is small, which corresponds to a perpendicular alignment. The Kolmogorov-Smirnov (KS) probability that these angular distributions are selected from an isotropic distribution is $\sim$ 4 $\times$ 10-5 for Milky Way-mass halos and $\sim$ 3 $\times$ 10-5 for cluster-mass halos, indicating that an isotropic, uncorrelated distribution of halo angular momenta and principle axes is excluded by the data. This alignment has previously been detected in works such as Shaw et al. (2006); Knebe et al. (2008); Deason et al. (2011); Hoffmann et al. (2014); Kiessling et al. (2015).
We can also quantify the alignment between host halo angular momentum vector and host major axis with the median angular separation between the two. For Milky Way-mass halos, the median angular separation is $\theta$ = $77\pm 3.0^{\circ}(17^{\circ})$ or $|\cos(\theta)|$ = 0.22 $\pm$ 0.05 (0.24). The first set of error values are the errors on the medians while the values in the parentheses correspond to the dispersions in the distribution of angles. For Cluster-mass halos, the median angular separation is $\theta$ = $75\pm 2^{\circ}(16^{\circ})$ or cos($\theta$) = 0.25 $\pm$ 0.04 (0.24). Errors are calculated by bootstrapping. For each set of angles, we sample the angles with replacement to create a mock sample of the same size and calculate its median. To get the uncertainty we repeat this 1000 times and calculate the standard deviation of the set of 1000 medians. For an isotropic distribution of angles, we would expect a median of 60, however, we are measuring angles that are approximately 15 greater. Angles significantly greater than 60 are indicative of perpendicular alignment which is consistent with the discussion in the previous paragraph that a halo’s angular momentum vector is preferentially perpendicular to its major axis.
Figure 3: The cumulative fraction of halos with angle less than $\rm cos(\theta)$ between the host halo angular momentum vector, $\vec{J}_{\mathrm{Host}}$ and its major axis $\vec{A}_{\mathrm{Host}}$ as a function of $\rm cos(\theta)$. The solid line is the result from the simulation data, the dashed line and the filled area are the CDF for an isotropic distribution and the area within the 14th and 86th percentiles. The percentiles are computed from random samples that are the same size as the simulation samples. Results for Milky Way-mass halos are shown in (a) and cluster-mass in (b).
5 Results II: Spatial Distribution of Subhalos
Next, we study the spatial distribution of subhalos about their hosts in two ways. First, we measure the degree to which the spatial positions of subhalos are aligned with the major axis of their host halo and whether or not subhalos reside in a planar structure roughly orthogonal to the host angular momentum. Second, we measure the way in which the 2d projection of the subhalo population belonging to a host changes when the halo is viewed from different directions. We quantify this projection effect primarily in terms of the projected mass fraction in subhalos.
5.1 Principle Axis and Angular Momentum Alignments of Subhalos and Their Hosts
We first discuss alignments between the major axis of the distribution of all subhalos belonging to a particular host halo and the major axis of said host halo. The principal axes of the subhalo component are defined as the eigenvectors of the inertia tensor corresponding to subhalo particles only. We calculated the angular separation between the major axis of the host, $\vec{A}_{\mathrm{Host}}$, and that of the subhalos, $\vec{A}_{\mathrm{Sub}}$. Previous works have shown that this result can vary as a function of the radius at which the calculation is performed (Schneider et al., 2012b). In this work, we chose to use major axis as defined up to the full virial radius of the halo. In Fig. 4, we show the cumulative fraction of halos with angle $|\rm cos(\theta)|$ between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ as a function of $|\rm cos(\theta)|$. The left plot (4(a)) shows results for Milky Way-mass halos and the right (4(b)) for cluster-mass halos.
In both panels of Fig. 4 we see that there is a deficit of halos at smaller values of $|\rm cos(\theta)|$, indicating that the angular separation between the major axis of the host halo matter distribution and the major axis of the subhalo distribution is smaller than it would be given an isotropic distribution. This implies, perhaps unsurprisingly, that $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ are preferentially aligned. In other words, the distribution of subhalos is aligned with the distribution of host halo mass. The KS probability that the result for Milky Way-mass halos (4(a)) is selected from an isotropic distribution is $P_{\rm KS}$ is $\sim 3\times 10^{-\rm 9}$. For cluster-mass halos (4(b)) this probability is $\sim 7\times 10^{-10}$.
For Milky Way-mass halos, the median angle between the host halo major axis the subhalo component major axis is $\theta$ = $30\pm 4.0^{\circ}(22^{\circ})$ or $|\cos(\theta)|$ = 0.86 $\pm$ 0.04 (0.26). For Cluster-mass halos, this median angle is $\theta$ = $16\pm 3^{\circ}(23^{\circ})$ or $|\cos(\theta)|$ = 0.96 $\pm$ 0.01 (0.24). Errors are calculated with the same bootstrap method as Section 4. For an isotropic distribution of angles (spanning from 0 to 90), we would expect a median of 60; however, we measure angular separations that are approximately 30- 45 smaller. This is consistent with the discussion in the previous paragraph that subhalos preferentially reside closer to the major axis of their host halo.
The alignment we find between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ is not unexpected, a fact to which we have already alluded. Within the model of hierarchical structure growth, subhalos merge with host halos in a correlated manner along filaments (e.g., Plionis & Basilakos, 2002; Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b; Shaw et al., 2006; Libeskind et al., 2007; Welker et al., 2014). In a recent work, Han et al. (2023) showed that $\sim 68\%$ of accreted subhalos enter their host halo through $\sim 38\%$ of the halo surface area defined at the virial radius. This results in alignment between the large-scale filamentary structure and that of a host halo (e.g., Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b). In turn, this causes the shape and orientation of halos to be ellipsoids with their longest axes pointed along the direction of the filament. Consequently, subhalos are preferentially located near this longest axis, the major axis, and this is the alignment that we detect.
Comparing 4(a) with 4(b), it is evident that the alignment between $\vec{A}_{\mathrm{Host}}$ and $\vec{A}_{\mathrm{Sub}}$ is more pronounced in cluster-mass halos than in Milky Way-mass halos. This difference may be attributed to the later formation times and larger subhalo abundances of cluster-mass halos, which generally have experienced significantly more recent mergers (Lacey & Cole, 1994; Wechsler et al., 2002; Zentner et al., 2005a). Consequently, this means that the major axes of cluster-mass halos have been more recently influenced by the influx of mergers along neighboring filaments. Collectively, these effects lead to subhalos being positioned more closely to the major axis of cluster-mass halos. Such a mass dependence on the strength of alignments has been proposed before for neighboring halos in works such as Jing (2002); Li et al. (2013); Schneider et al. (2012a); Xia et al. (2017). A follow-up analysis would be necessary in order to confirm this interpretation.
In principle, it may be possible for the major axes of host halo and subhalo components to be aligned without an anisotropic distribution of subhalos (though this seems inconsistent with our other results, see below). Suppose that the distribution of subhalos is characterized by principle axis ratios $(b/a)_{\mathrm{subs}}$ and $(c/a)_{\mathrm{subs}}$ (from the subhalo only samples) that are significantly larger than the principle axis ratios determined by the overall mass of the host halos (from the host only) samples. For example, if $(c/a)_{\mathrm{subs}}\sim 1\gg(c/a)$ and $(b/a)_{\mathrm{subs}}\sim 1\gg(b/a)$, then the subhalo mass would be distributed nearly isotropically, so that axis alignments would be far less noteworthy. However, this is not the case. We find that subhalo mass is distributed approximately as anisotropically as its host halo’s mass as quantified by axis ratios. The axis ratios for host halos in the Milky Way-mass simulations are $(b/a)=0.86\pm 0.10$ and $(c/a)=0.74\pm 0.11$, while the corresponding ratios for subhalo mass are $(b/a)_{\mathrm{subs}}=0.84\pm 0.16$ and $(c/a)_{\mathrm{subs}}=0.66\pm 0.16$. Similarly, for cluster mass halos the axis ratios are $(b/a)=0.78\pm 0.08$ and $(c/a)=0.68\pm 0.07$ for host halos while they are $(b/a)_{\mathrm{subs}}=0.60\pm 0.2$ and $(c/a)_{\mathrm{subs}}=0.36\pm 0.18$ for the subhalos. Subhalo mass is distributed at least as anisotropically as the host halo mass indicating that alignments among principle axes designate meaningful alignments of a strongly anisotropic distribution of subhalos. In fact, our results suggest that the distribution of subhalo mass is somewhat less spherical than the overall mass distribution of a halo.
Figure 4: The cumulative fraction of halos with angle $|\rm cos(\theta)|$ between the host halo major axis $\vec{A}_{\mathrm{Host}}$ and that of its subhalos, $\vec{A}_{\mathrm{Sub}}$, as a function of $|\rm cos(\theta)|$. The solid line is the result from the simulation data, the dashed line and the filled area are the CDF for an isotropic distribution and the area within the 14th and 86th percentiles. Panel (a) is for Milky Way mass halos and (b) is for cluster mass halos.
In Fig. 5, we compare the angular momentum vectors of host halos, $\vec{J}_{\mathrm{Host}}$, to that of their subhalos, $\vec{J}_{\mathrm{Sub}}$. $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$ are both defined in Section 3.2. Notice that, unlike orientations relative to the major axis, the x-axis in Fig. 5 is $\rm cos(\theta)$ rather than $|\rm cos(\theta)|$ to display cases in which subhalo angular momenta are anti-parallel to that of their host. The left panels correspond to Milky Way-mass halos and the right to cluster-mass halos. In the top row, figures $a$ and $b$, we see a similar trend as in Fig. 4 in which there is a deficit of halos at smaller values of $\rm cos(\theta)$, indicating that there is alignment between $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$. The KS probability that the distribution for Milky Way-mass halos in the top panel of 5(a) is selected from an isotropic distribution is $P_{\rm KS}$ $\sim 1\times 10^{-\rm 9}$. For cluster-mass halos in the top panel of 5(b), $P_{\rm KS}$ is $\sim 4\times 10^{-\rm 5}$. As we discussed in the preceding paragraph, the tendency for subhalo and host halo angular momenta to be aligned makes sense within the context of hierarchical growth – subhalos merging in from coherent directions along filaments will preferentially orbit near a plane containing the major axis. The fact that we detect alignment between major axes strengthens this argument.
The other two rows of panels in Fig. 5 show our results for alternative methods of measuring the subhalo angular momentum. In the first case, we define the angular momentum of each subhalo using the position and velocity of its center as specified in the halo catalog and sum over that angular momentum for all subhalos within a host to find the total angular momenta of the subhalo component of the halo. We will refer to this measure of angular momentum as $\vec{J}_{\mathrm{peak}}$. This method is distinct from the previous method in which the subhalo angular momenta were computed by summing over the particle content of each subhalo. In panels $c$ and $d$ we show the cumulative distribution of angles between $\vec{J}_{\mathrm{peak}}$ and $\vec{J}_{\mathrm{Host}}$. The distributions in $c$ and $d$ are, unsurprisingly, both similar to those of $a$ and $b$. This suggests that our alignment results are insensitive to the details of the method used to determine subhalo angular momenta. In the second alternative case, we display the alignments of angular momenta for individual subhalos, rather than for populations of subhalos associated with individual hosts. In particular, we compute the angle between the angular momentum of each individual subhalo and the angular momentum vector of its host. To distinguish this measurement, we refer to the angular momentum of a single subhalo as $\vec{J}_{\mathrm{ind}}$. We sum over the angular momenta of all particles in any individual subhalo to determine $\vec{J}_{\mathrm{ind}}$. The result of this exercise is shown in the bottom panels, $e$ and $f$. The resulting distribution of alignments differs from those in panels $a$, $b$, $c$, and $d$, because in this case the individual subhalos corresponding to a single host can fall into distinct bins of zenith angle, whereas in the previous calculation each host had a single zenith angle corresponding to the alignment of the host angular momentum with the net angular momentum of the population of subhalos. This has the effect of spreading the contribution from the subhalo population associated with any particular host galaxy among a range of zenith angle values. Furthermore, the alignment distribution computed in this way is significantly less noisy because the distribution is assembled from many more data points. Calculating angles for individual halos allows us to decipher whether or not the alignment in the top two figures, $a$ and $b$, is due to a large fraction of subhalos having angular momenta aligned with that of their hosts, or, subhalo angular momenta that are spread out but average to a value close to that of its host. The anisotropy is weaker for $\vec{J}_{\mathrm{ind}}$, implying that those subhalos with the largest angular momenta are more well-aligned with the host angular momentum.
Finally, we quantify the alignment between angular momenta with the median angular separation between host and subhalo angular momentum vectors. For Milky Way-mass halos, the median angular separation is $\theta$ = $50\pm 9^{\circ}(38.0^{\circ})$ or cos($\theta$) = 0.63 $\pm$ 0.12 (0.55). For Cluster-mass halos, it is $\theta$ = $56\pm 5.5^{\circ}(34^{\circ})$ or cos($\theta$) = 0.56 $\pm$ 0.08 (0.49). Errors are calculated by bootstrapping using the same method as in Section 4. For an isotropic distribution of angles (spanning from 0 to 180), we would expect a median of 90, with 0 indicative of perfect alignment between the angular momentum vectors, and 180 indicative of $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$ pointing in complete opposite directions. Our findings show that the median angle of separation between angular momentum of the host and subhalos is significantly smaller than 90, indicative of strong alignment between the two, and is consistent with the discussion in the previous paragraph that subhalos should orbit in the same plane as the host halo.
Figure 5: The same as Fig. 4 except for the angles between the host and subhalo angular momentum vectors, $\vec{J}_{\mathrm{Host}}$ and $\vec{J}_{\mathrm{Sub}}$, respectively. In the bottom 4 figures, we use alternative measurements of subhalo angular momentum. For comparison with the original method of measuring subhalo angular momentum, we also plot the lines from the top plots in a lighter shade along side the alternative method result. In the middle row, subhalo angular momenta is defined by the position and velocity of its density peak. Like in the top row figures, we sum over all subhalo angular momenta to define $\vec{J}_{\mathrm{peak}}$. In the bottom row, $\vec{J}_{\mathrm{ind}}$ corresponds to the angular momentum of a single subhalo found by summing over all its particles rather than the sum of particles over all subhalos belonging to a host.
Although we anticipate that the alignments of $\vec{A}_{\mathrm{Sub}}$ and $\vec{A}_{\mathrm{Host}}$, and, $\vec{J}_{\mathrm{Sub}}$ and $\vec{J}_{\mathrm{Host}}$ are a result of mergers along coherent directions, such as filaments, we should expect significant scatter in these alignments. Cluster-sized halos are comparable in size to the filament in which they reside and may also be located at a node connecting multiple filaments. Galaxy-sized halos at present are often smaller than the size of the filaments in which they live but may have accreted their surviving subhalos at times when filaments were thinner (Vera-Ciro et al., 2011). The distribution of subhalos that merge from other directions or from presently-subdominant filamentary structure introduces considerable scatter into the anisotropic distributions we have found. Furthermore, simulations have shown that spin alignment with filaments has a mass dependence (Bailin & Steinmetz, 2005; Aragón-Calvo et al., 2007; Paz et al., 2008; Zhang et al., 2009; Codis et al., 2012; Libeskind et al., 2011, 2015; Forero-Romero et al., 2014). In particular, Codis et al. (2012) find that the orientation of a halo’s spin relative to the filament in which it resides has a mass dependence with a turnover from pointing parallel to the filament to perpendicular at a critical mass of $\sim$ 5($\pm$1) $\times$ $10^{12}\text{M}_{\odot}$. This mass is greater than that of our Milky way-mass halos and considerably lower than that of our cluster-mass halos. This suggests that Milky Way-mass halos will be oriented differently relative to the present-day large-scale structure than cluster-mass halos will. This is interesting to note and may be important for a variety of forthcoming applications, including gravitational lensing studies, but we will not pursue this further in this work.
5.2 The Distribution of Subhalo Mass as a Function of Viewing Angle
In order to interpret strong gravitational lensing measurements, it is most practical to focus on the mass in the central region of the halo and specifically, the mass within approximately the Einstein radius in projection. For lenses at z=0.5 composed of a galaxy within a dark matter halo within the mass range of our Milky Way-mass halos, with a source at z=2, we estimate an Einstein radius of approximately 2 arcsecs or 12.2 kpc and for lenses composed of a galaxy in a cluster-mass dark matter halo they are approximately 10 arcsecs or 61 kpc. Because of the ellipsoidal shape of halos, the projected matter density near the center will vary depending on the orientation from which the halo is viewed and a greater projected matter density will be observed near the center when the halo is viewed along its long axis compared to its shortest. We predict that a greater projected density of subhalos at the halo center will also be observed when the halo is viewed along this axis.
Determining the influence of subhalos on strong lensing systems is a topic of great contemporary interest. We have already demonstrated that subhalos are distributed anisotropically around their hosts and that subhalos are found preferentially aligned the long axes of their hosts. We quantify this effect to inform its importance in previous and forthcoming lensing measurements. We show the results from this analysis in Fig. 6 where the left panels correspond to Milky Way-mass halos and the right panels to cluster-mass halos. The panels of Fig. 6 depict the mass fraction in subhalos in a cylindrical projection as a function of both radius of the cylindrical projection (top panels) and viewing angle relative to the major axis of the halo (bottom panels). The cylinder scheme was discussed in Section 3 and it may be useful to review Fig. 2 to visualize the geometry under consideration. In all panels, the vertical axis is the ratio of the projected subhalo mass fraction enclosed within a particular cylinder, oriented at an angle $\theta$ away from the halo major axis, divided by the median mass fraction for all cylindrical projections across all halos and angles.
Figure 6: The top two panels show subhalo mass fraction captured within cylinders projected along predefined axes on a uniformly spaced grid on the surface of a sphere normalized by the median mass fraction across all halos and angles. The left panel in blue corresponds to Milky Way-mass halos and the right panel in pink to cluster-mass halos. The x-axis is the ratio of the radius of the cylinder to virial radius of the halo, and, the y-axis is the normalized subhalo mass fraction. The solid lines are the median normalized mass fraction within bins of $r/r_{\rm{VIR}}$ at four different angular separations ranging from aligned with the major axis (lightest shaded line) to perpendicular to the major axis (darkest shaded line). The bottom two panels contain the same information presented in a different way. For these figures, the x-axis is the angular separation. The solid and dotted lines correspond to the mass fraction within cylinders of radii equal to approximately 10% and 5% of the virial radius. The shaded regions are the error margins corresponding to 14th and 86th percentiles.
In the top two panels of Fig. 6, the solid line is the median normalized subhalo mass fraction within bins of $r/r_{\rm{VIR}}$ at four different angular separations ranging from aligned with the major axis (lightest shaded line) to perpendicular to the major axis (darkest shaded line). The shaded regions are the error margins corresponding to 14th and 86th percentiles. Mass fractions are normalized by the median subhalo mass fraction across all halos and angles. Consequently, if the subhalos were distributed isotropically, the solid lines in each panel would be horizontal lines with a value of 1.0. In both Milky Way-mass and cluster-mass halos, the subhalo mass fraction increases as cylinder orientation approaches the major axis. Smaller cylinder radii lead to dramatically larger differences, which makes sense because smaller cylinder radii correspond to averaging over a smaller physical scale so that correlations are not washed out by subhalos that are poorly aligned with the major axis, but fall within the cylinder radius. This effect is more dramatic for cluster-mass halos where the mass fraction is approximately 1.95 times greater at small cylinder radii along the major axis compared to perpendicular, whereas the trend for Milky Way-mass halos is both smaller (1.75 times greater at small radii along the major axis) and noisier (due to the smaller sample size). The bottom two panels show the median normalized mass fraction within bins of separation angle at two cylinder radii equal to approximately 5% and 10% of the virial radius. Once again, the shaded regions are the bootstrapped error margins corresponding to 14th and 86th percentiles. In these figures we again note an increase in mass fraction as cylinder orientation approaches the major axis for the same reason as discussed above. We note that the difference in maximum mass fraction between the top and bottom figures is due to the fact that in the bottom figures we are binning data in cos($\theta$).
This result naturally emerges from the anisotropies already discussed in Section 5.1. We saw overall stronger alignments with particular axes for cluster-mass halos in Section 5.1 as well, so the difference in the strength of this effect between Milky Way-mass halos and cluster-mass halos is not surprising. We attribute the noisiness in the Milky Way result to the fact that there are fewer than half as many Milky Way-mass host halos compared to cluster-mass host halos. Moreover, smaller host halos tend to have a smaller total fraction of their mass in subhalos, so the Milky Way-mass hosts contain fewer subhalos and thus there is a larger shot noise associated with subhalo position.
For lensing, a massive subhalo of some mass, $M$, in the line of sight will leave a larger lensing signal than a collection of smaller subhalos with collective mass, $M$. To this end, we ran a smaller scale version of the analysis discussed above with a variety of subsamples computed using threshold samples of subhalos above a minimum mass. We calculate projected subhalo mass for subhalos of mass, $M$, greater than 0.025, 0.05, 0.075, and 0.1 times the total halo virial mass. These numbers are selected such that they fall within the range of available subhalo masses – Milky Way-mass halos in our sample do not have subhalos more massive than 0.2 times their total halo virial mass and only a handful of cluster-mass halos have subhalos that exceed this threshold. For this iteration of the analysis, we calculate the mass fraction at only two radii on the scale of the expected Einstein radius, 0.05 and 0.1 times the halo virial radius, along 500 randomly selected axes through the halo. We find that as this lower mass limit is increased, there is a small increase in the overabundance of projected subhalo mass near the major axis, although, the change is not statistically significant. When considered together with our other results indicating subhalo anisotropies, we nevertheless encourage the reader to be mindful of this potential effect.
6 Conclusion and Discussion
6.1 Conclusion
In this paper we have shown that subhalos are distributed anisotropically with respect to their host halos. In particular, subhalos tend to reside preferentially near the major axis of their host halo. We examined specifically the following relationships between the host halo and its subhalo population: (1) the degree to which the angular momentum of the host halo is correlated with the net angular momentum of its associated subhalo population; (2) the alignment of the principle axes of the host halo with the principle axes defined by its subhalo population; and (3) we measured the projected mass in subhalos as a function of projection angle relative to the major axis of the host. This study anticipates applying these results to the interpretation of strong lensing observations, but we have phrased the alignment in a variety of ways so that these alignment results may be applied more broadly in a variety of other applications. To our knowledge, this is the only analysis of this type performed using zoom-in simulations of comparably high resolution. This has enabled us to study subhalo-host relationships in greater detail than previous studies. Our main results are as follows.
  • Both cluster-mass and Milky Way-mass host halo’s major axes are generally well aligned with those of their subhalos. In this broad sense, subhalos are preferentially oriented in the same way as the overall mass distribution of their host halos. For Milky Way-mass halos, the median angle between the host halo major axis the subhalo component major axis is $\theta$ = $30\pm 4^{\circ}$ or $|\cos(\theta)|$ = 0.86 $\pm$ 0.04. For Cluster-mass halos, this median angle is $\theta$ = $25\pm 3.5^{\circ}$ or $|\cos(\theta)|$ = 0.91 $\pm$ 0.03.
  • Both cluster-mass and Milky Way-mass host halo’s angular momenta are fairly well aligned with the net angular momenta of their subhalos. For Milky Way-mass halos, the median angle between the host halo angular momentum the net subhalo angular momentum is $\theta$ = $50\pm 9^{\circ}$ or cos($\theta$) = 0.63 $\pm$ 0.12. For Cluster-mass halos, this median angle is $\theta$ = $56\pm 5.5^{\circ}$ or cos($\theta$) = 0.56 $\pm$ 0.08.
  • Halos that are viewed along their major axis have a greater projected subhalo mass near their centers compared to those viewed along a random axis direction. As seen in the top panels of Fig. 6, within projected radii of a few percent of the virial radius of the host halo, the fraction of mass in subhalos is $\sim 175\%$ larger for Milky Way mass halos and as much as $\sim 195\%$ larger for cluster halos when projected along the major axis as compared to the average from a random projection. To within our ability to measure, this preference does not have a strong dependence upon the mass of the subhalo.
6.2 Discussion
The intrinsic alignment between halo major axes provides insight into structure growth and halo formation. Within hierarchical formation models, overdense regions of the density field collapse into sheets and filaments and underdense regions into voids. Dark matter halos form at peaks of the density field which collapse and grow over time through mergers and smooth accretion.
Broadly speaking, the alignments that we have explored can be understood within the standard paradigm of hierarchical structure formation. Early-forming halos collapse from peaks in the initial density field and form triaxial structures (Zel’dovich, 1970). Tidal torquing then produces net halo angular momenta that are roughly orthogonal to halo major axes (Peebles, 1969; White, 1984). Those halos which reside within larger-scale overdensities then merge to form larger halos and so the population of massive halos is built from mergers with smaller halos. In the net, these mergers take place along strongly correlated directions (Zentner et al., 2005b; Libeskind et al., 2007, 2010). In particular, halos preferentially merge along the filamentary structure. Sequences of mergers from preferential directions cause halos to become elongated in the direction of the merger. As a result, host halos tend to be prolate and their major axes tend to point along the filaments (Jing & Suto, 2002; Bailin & Steinmetz, 2005; Faltenbacher et al., 2005; Zentner et al., 2005b). The very same mergers supply host halos with their subhalos and the subhalos tend to continue on elongated orbits that reflect the direction from which they merged (Zentner et al., 2005b). Of course, the details of these correlations must be computed from numerical simulations and many of the papers cited in this paragraph have done just that.
This overall picture is supported by observational studies. Halo ellipticity has now been detected via weak gravitational lensing and those lensing measurements also confirm that the major axes of galaxies must be correlated with the major axes of their host halos. The non-zero ellipticities of halos have been detected in both weak lensing (e.g., Evans & Bridle, 2009; Clampitt & Jain, 2016; Gonzalez et al., 2021; Robison et al., 2023) and strong lensing (e.g., Okabe et al., 2020). Moreover, these measurements are broadly consistent with theoretical expectations (e.g., Bett, 2012; Velliscig et al., 2015; Chisari et al., 2017). Of particular interest in the current context are the results of Evans & Bridle (2009) and Gonzalez et al. (2021) who found that the orientation of the dark matter ellipse was correlated with the distribution of member galaxies in clusters. This suggests a detection of the bias for finding subhalos to be aligned with the major axes of their host halos. There are also a number of studies of SDSS data that show that satellite galaxies tend to reside along the major axes of their host galaxies (Brainerd, 2005; Yang et al., 2006; Wang et al., 2018).
The alignments between host dark matter halos and their subhalos has implications for a variety of observations and particularly for the analysis of strong gravitational lensing. Strong gravitational lensing has become an indispensable probe of galaxy and halo structure. Among the most ambitious achievements and aims of strong lensing is to identify lensing features caused by subhalos within a primary lens system and thereby constrain the properties of the dark matter. This can be done using flux-ratio anomalies (Mao & Schneider, 1998; Metcalf & Madau, 2001; Dalal & Kochanek, 2002; Bradač et al., 2002; Keeton, 2003; Metcalf & Zhao, 2002; Metcalf et al., 2004; Treu & Marshall, 2016; Gilman et al., 2018; Harvey et al., 2019; Treu et al., 2022; Zelko et al., 2024; Nierenberg et al., 2024) and perturbations in lensing arcs (Meneghetti et al., 2006; Vegetti et al., 2010; Despali et al., 2018; Minor et al., 2021). In each case, it is necessary to know the amount of substructure one should expect near the Einstein radius of the lens systems as viewed in projection along the line of sight.
We found that the number of subhalos as well as the projected mass fraction in subhalos is greater when viewed along a two-dimensional projection aligned with the major axis of the host halo. The mass fraction in subhalos declines as the viewing angle with respect to the principle axis increases. The halos themselves are typically nearly prolate ellipsoids, so the total mass density in projection is higher when viewed along a direction closely aligned with the major axis. This suggests that halos may be more likely to serve as strong lenses when viewed at an angle nearly aligned with the major axis, though the strength of this preference depends upon the poorly constrained degree of alignment between the principle axes of a lens halo and the principle axes of the galaxy it hosts. Indeed, Hennawi et al. (2007) estimate that the mean value of the cosine of the viewing angle with respect to the major axis, $|\cos(\theta)|$, should be $\sim 0.67$ for multiply-imaged systems, rather than $0.5$, as would be expected if lensing probability were independent of viewing angle. There is a clear implication of these results. If halos that serve as strong lenses are observed in projections that are nearly projections along the major axis, at least preferentially, then the expected mass fraction in subhalos should be substantially larger than predictions for the global mass fraction in subhalos. In follow-up work, we will explore the degree to which this potential bias can affect the interpretation of strong lensing observations and we advise those analyzing strong lensing data to be aware of this potential bias.
There are possibly other important consequences of the alignments highlighted in this work. Our finding that subhalo and host halo angular momenta are aligned has consequences for galaxy formation studies. For example, it is common practice for semi-analytic galaxy formation models to adopt host halo angular momentum as a proxy for galaxy angular momentum either probabilistically (Somerville et al., 2008; Guo et al., 2011) or deterministically (Benson, 2012). Interestingly, our results suggest that a large fraction of the net angular momentum of the host halo is carried by the subhalos reside on orbits that are close to the plane of any galactic disk. Sufficiently large subhalos will likely host their own galaxies. This provides an interesting contrast to a now well-known observation in the Local Group. The dwarf satellite galaxies of the Milky Way seem to orbit the Milky Way in a plane that is nearly perpendicular to the galaxy disk so that the angular momentum carried by the dwarf satellites is nearly perpendicular to that angular momentum of the Milky Way galaxy (Kroupa et al., 2005; Pawlowski, 2018). Satellites of Andromeda are found to be distributed anisotropically (Doliva-Dolinsky et al., 2023) and orbit the disk of Andromeda in a separate plane at a tilt of $\sim$50 (Pawlowski, 2018) with respect to the angular momentum of Andromeda. Such planar orbits of satellites have also been found in simulations using Milky Way and Andromeda galaxy analogs in Santos-Santos et al. (2020) and Samuel et al. (2021) with the latter showing that they can exist in a plane perpendicular to the galactic disk. However, recent work indicates that the perpendicular alignment seen in the Milky Way could be temporary and due to a chance alignment with its two most distant satellites (Leo I and II) (Sales & Navarro, 2023). Additionally, prior simulation work by Fielder et al. (2018) found host halo angular momentum to be the weakest indicator of subhalo abundance, further reinforcing the idea of alignment between host halo and subhalo angular momenta. All things considered, we believe that caution must be taken by galaxy formation studies when defining galaxy angular momentum.
In this paper, we provide a comprehensive look at the biased distribution of subhalos about their host halos. Using zoom-in simulations, we have shown that subhalos are distributed anisotropically with respect to their host halos. We find that (i) they are preferentially located along the major axis of their host halo and (ii) they orbit in a plane perpendicular to the spin axis of their host halo. These are effects that may be potentially important in galaxy formation models and, in particular, in the interpretation of strong lens systems.
Acknowledgements
We thank Atınç Çağan Şengül for insightful comments and helpful discussion that improved this manuscript.
This research made use of Python, along with many community-developed or maintained software packages, including IPython (Pérez & Granger, 2007), Jupyter (jupyter.org), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), Pandas (McKinney, 2010), and SciPy (Jones et al., 2001). This research made use of NASA’s Astrophysics Data System for bibliographic information.
This work used data from the Symphony suite of simulations (http://web.stanford.edu/group/gfc/symphony/), which was supported by the Kavli Institute for Particle Astrophysics and Cosmology at Stanford University and SLAC National Accelerator Laboratory, and by the U.S. Department of Energy under contract number DE-AC02-76SF00515 to SLAC National Accelerator Laboratory.
Data Availability
The Symphony data products used in this work are publicly available at phil-mansfield.github.io/symphony.